Random samples of size 3 are taken from a population of the numbers 3, 4,5,6, 7,8,and 9.


1. How many samples are possible? List them and compute


the mean of each sample. 2. Construct the sampling distribution of the sample means.


3. Construct the histogram of the sampling distribution of


the sample means. Describe the shape of the histogram.


1
Expert's answer
2022-06-14T00:16:39-0400

We have population values 3, 4, 5, 6, 7, 8, 9, population size N=7 and sample size n=3.

Mean of population "(\mu)" = "\dfrac{3+4+5+6+7+8+9}{7}=6"

Variance of population 


"\sigma^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n}""=\dfrac{1}{7}(9+4+1+0+1+4+6)=4""\sigma=\sqrt{\sigma^2}=\sqrt{4}=2"

1. Select a random sample of size 4 without replacement. We have a sample distribution of sample mean.

The number of possible samples which can be drawn without replacement is "^{N}C_n=^{7}C_3=35."

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 no & Sample & Sample \\
& & mean\ (\bar{x})
\\ \hline
 1 & 3,4,5 & 12/3 \\
 \hdashline
 2 & 3,4,6 & 13/3 \\
 \hdashline
 3 & 3,4,7 & 14/3\\
 \hdashline
 4 & 3,4,8 & 15/3 \\
 \hdashline
 5 & 3,4,9 & 16/3 \\
 \hdashline
 6 & 3,5,6 & 14/3 \\
 \hdashline
 7 & 3,5,7 & 15/3\\
 \hdashline
 8 & 3,5,8 & 16/3 \\
 \hdashline
 9 & 3,5,9 & 17/3\\
 \hdashline
 10 & 3, 6,7 & 16/3 \\
 \hdashline
 11 & 3,6,8 & 17/3 \\
 \hdashline
 12 & 3,6,9 & 18/3 \\
 \hdashline
 13 & 3,7,8 & 18/3 \\
 \hdashline
 14 & 3,7,9 & 19/3 \\
 \hdashline
 15 & 3,8,9 & 20/3 \\
 \hdashline
 16 & 4,5,6 & 15/3 \\
 \hdashline
 17 & 4,5,7 & 16/3 \\
 \hdashline
 18 & 4,5,8 & 17/3 \\
 \hdashline
 19 & 4,5,9 & 18/3 \\
 \hdashline
 20 & 4,6,7 & 17/3 \\
 \hdashline
 21 & 4,6,8 & 18/3 \\
 \hdashline
 22 & 4,6,9 & 19/3 \\
 \hdashline
 23 & 4,7,8 & 19/3 \\
 \hdashline
 24 & 4,7,9 & 20/3 \\
 \hdashline
 25 & 4,8,9 & 21/3 \\
 \hdashline
 26 & 5,6,7 & 18/3 \\
 \hdashline
 27 & 5,6,8 & 19/3 \\
 \hdashline
 28 & 5,6,9 & 20/3 \\
 \hdashline
 29 & 5,7,8 & 20/3 \\
 \hdashline
 30 & 5,7,9 & 21/3 \\
 \hdashline
 31 & 5,8,9 & 22/3 \\
 \hdashline
 32 & 6,7,8 & 21/3 \\
 \hdashline
 33 & 6,7,9 & 22/3 \\
 \hdashline
 34 & 6,8,9 & 23/3 \\
 \hdashline
 35 & 7,8,9 & 24/3 \\
 \hdashline
\end{array}"



2.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f(\bar{X}) &\bar{X} f(\bar{X}) & \bar{X}^2f(\bar{X})
\\ \hline
 12/3 & 1/35 & 12/105 & 144/315 \\
 \hdashline
13/3 & 1/35 & 13/105 & 169/315 \\
 \hdashline
 14/3& 2/35 & 28/105 & 392/315 \\
 \hdashline
 15/3 & 3/35 & 45/105 & 675/315 \\
 \hdashline
 16/3 & 4/35 & 64/105 & 1024/315\\
 \hdashline
 17/3 & 4/35 & 68/105 & 1156/315 \\
 \hdashline
 18/3 & 4/35 & 90/105 & 1620/315 \\
 \hdashline
 19/3 & 4/35 & 76/105 & 1444/315 \\
 \hdashline
 20/3 & 4/35 & 80/105 & 1600/315 \\
 \hdashline
 21/3 & 3/35 & 63/105 & 1323/315 \\
 \hdashline
 22/3 & 2/35 & 44/105 & 968/315 \\
 \hdashline
 23/3 & 1/35 & 23/105 & 529/315 \\
 \hdashline
 24/3 & 1/35 & 24/105 & 576/315 \\
 \hdashline
\end{array}"



Mean of sampling distribution 


"\mu_{\bar{X}}=E(\bar{X})=\sum\bar{X}_if(\bar{X}_i)=\dfrac{630}{105}=6=\mu"



The variance of sampling distribution 


"Var(\bar{X})=\sigma^2_{\bar{X}}=\sum\bar{X}_i^2f(\bar{X}_i)-\big[\sum\bar{X}_if(\bar{X}_i)\big]^2""=\dfrac{11620}{315}-(6)^2=\dfrac{8}{9}= \dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})"




"\sigma_{\bar{X}}=\sqrt{\dfrac{8}{9}}=\dfrac{2\sqrt{2}}{3}\approx0.9428"


3.



Symmetric distribution.